two automobiles set out simultaneously from places 100 miles apart. they traveled toward each other and passed at the end of 1 hour and 40 minutes. the speed of one automobile was 5 miles per hour greater than of the other. find the speed of each. can you please answer it for me? thanks!
Answer by mananth(16949)

Distance they meet from one car = x
other car travels 100-x miles when they meet.
distance/time = speed
1 h 40 mins = 1 2/3 hours = 5/3 hours

x/(5/3) -(100-x)/(5/3)= 5
3x/5 - 3(100-x)/5 =5
(3x-300+3x)/5 = 5
6x-300=25
6x= 325
x=54.16
one car travels 54.16 miles in 5/3 hours
so in 1 hour it travels 54.16/(5/3)
= 32.5 mph
second car travels 100-54.16 = 45.84 mile
speed = 45.84/(5/3)= 27.5 mph
cheers